OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y =1/3x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (lalaly):

y=mx+C if its perpendicular then the slope is -3 y=-3x+c plug in (4,-2) -2=-3(4)+c c=-2+12 = 10 so the equation is y=-3x+10